# UniDoc-Bench: A Unified Benchmark for Document-Centric Multimodal RAG

### v1

@@ Abstract
Multimodal retrieval-augmented Generation (MM-RAG) is a key approach for applying large language models (LLMs) and agents to real-world knowledge bases, yet current evaluations are fragmented—focusing on either text or images in isolation, or simplified multimodal setup, failing to capture document-centric multimodal use cases. In this paper, we introduce UniDoc-Bench [^1] The code and data will be available at: https://github.com/SalesforceAIResearch/UniDOC-Bench , the first large-scale, realistic benchmark for MM-RAG built from 7070k real-world PDF pages across $8$ domains. Our pipeline extracts and links evidence from text, tables, and figures, then generates $1,600$ multimodal QA pairs spanning factual retrieval, comparison, summarization, and logical reasoning queries. To ensure reliability, $20\%$ of QA pairs are validated by multiple annotators and expert adjudication. UniDoc-Bench supports apples-to-apples comparison across four paradigms — 1) text-only, 2) image-only, 3) multimodal text–image fusion and 4) multimodal joint retrieval — under a unified protocol with standardized candidate pools, prompts, and evaluation metrics. Our experiments show that multimodal text–image fusion RAG systems consistently outperform both unimodal and jointly multimodal embedding–based retrieval, indicating that neither text nor images alone are sufficient and that current multimodal embeddings remain inadequate. Beyond benchmarking, our analysis reveals when and how visual context complements textual evidence, uncovers systematic failure modes, and offers actionable guidance for developing more robust MM-RAG pipelines.
@@ 1 Introduction
Retrieval-augmented generation (RAG) has become a widely used approach for applying large language models (LLMs) and agents to real-world knowledge bases (Gao et al., 2023; Fan et al., 2024). The dominant text-only pipeline applies Optical Character Recognition (OCR) (Li et al., 2022; Xue et al., 2024; Poznanski et al., 2025) to flatten document pages into text, indexes them as chunks, retrieves top-k text passages, and feeds them to a generator. However, many answers depend on information embedded in figures, charts, tables, and complex layouts, where OCR often discards crucial spatial and visual semantics (e.g., map, axes, bar lengths, color encodings) (Ma et al., 2024a; Faysse et al., 2024a). These limitations have driven the rapid development of multimodal RAG (MM-RAG), which embeds documents across modalities (text, tables, and images) and retrieves and reasons over them jointly, emerging as a key paradigm for document intelligence.
Current MM-RAG evaluation benchmarks exhibit substantial limitations, as summarized in Table 1. Many are restricted to a single image or a single document page as reference (Mathew et al., 2021; 2022; Zhu et al., 2022; Li et al., 2024; Ma et al., 2024b), cover narrow domains Mathew et al. (2021; 2022); Zhu et al. (2022); Li et al. (2024), under-represent modalities (Li et al., 2024; Mathew et al., 2022), operate at limited scale (few queries/pages)  (Ma et al., 2024b; Wang et al., 2025b) or lack a highly relevant database for RAG evaluation (Ma et al., 2024b). These gaps hinder fair and comprehensive comparison across methods. Moreover, debatable claims have emerged — such as that “image retrieval is all you need” (Faysse et al., 2024a; Su et al., 2025) or that multimodal retrieval is inherently superior (Zhang et al., 2024b; Yu et al., 2024b)— without enough fair and unified evaluation. In response, we introduce UniDoc-Bench, a manually verified benchmark spanning $8$ domains and covering text, chart, and table content, explicitly designed for cross-modality grounding with examples shown in Figure 1. Crucially, UniDoc-Bench enables apples-to-apples evaluation of text-retrieval, image-retrieval, multimodal text-image-fusion retrieval, and multimodal joint retrieval pipelines using highly relevant large document database and multi-type, cross-modality-grounding queries under a unified protocol. This setup provides an unbiased view of when multimodal retrieval offers advantages beyond single modalities. In practice, UniDoc-Bench quantifies multimodal gains, guides system design choices, and accelerates the development of effective MM-RAG systems for real-world document intelligence.
@@
We curate a high-quality multimodal RAG evaluation benchmark by designing and applying a classification-based filtering scheme to unlabeled, real-world PDF documents (PDFA (Montalvo & Wightman, 2024)), yielding 7070k highly relevant pages across eight widely used domains —Finance, Legal, Healthcare, Commerce and Manufacturing, CRM, Energy, Education, and Construction—containing rich cross-modality content, including text, tables, and images. We construct a knowledge graph that links cross-modality contents across documents via overlapping entities, and leverage these connections to synthesize 1,600 QA pairs spanning four question types: factual retrieval, comparison, summarization, and logical reasoning, enabling multi-modality grounding and reflecting realistic retrieval scenarios. To ensure quality, $20\%$ of the QA pairs are evaluated by three independent annotators for faithfulness, completeness, self-containment, human intent, and evidence usability, with disagreements resolved through expert adjudication. Figure 2 illustrates the full pipeline from PDF segmentation to dataset creation and evaluation.
@@
Table 1: Comparison of existing dataset with UniDoc-Bench.
Benchmarks Domain Evidence # Queries # Pages RAG Unified Multiple
of Doc Suitable Evaluation Reference
ArxivQA (Li et al., 2024) single [Uncaptioned image] 0.5k -
TAT-DQA (Zhu et al., 2022) single [Uncaptioned image]  [Uncaptioned image] 1.6k -
InfoVQA (Mathew et al., 2022) single [Uncaptioned image] 0.5k -
DocVQA (Mathew et al., 2021) single [Uncaptioned image]  [Uncaptioned image] 0.5k -
MMLONG (Ma et al., 2024b) multiple [Uncaptioned image]  [Uncaptioned image][Uncaptioned image] 1k 6k
REALMM (Wasserman et al., 2025) multiple [Uncaptioned image]  [Uncaptioned image][Uncaptioned image] 5k 8k
ViDoSeek (Wang et al., 2025b) multiple [Uncaptioned image]  [Uncaptioned image][Uncaptioned image] 1.2k 10k
@@
  • Unified Evaluation: Apples-to-apples comparison across different baseline RAG systems. Multiple Reference: Supports multi-hop, multi-modality, multi-source grounding.
In this paper, we make the following contributions:
  • We introduce a new multimodal RAG benchmark built from real-world PDF documents, comprising 70k pages across 8 domains, with 1,600 human-verified QA pairs referencing text, figures, and tables, spanning $4$ question types.
  • We present an associated data synthesizing pipeline for creating multimodal RAG evaluation datasets, designed to be compatible with any document database.
  • We propose a fair and reproducible evaluation framework by fixing candidate pools across modalities , and measuring retrieval effectiveness, answer faithfulness, and completeness end-to-end across different RAG systems. Specifically, to ensure fairness when comparing against text-only RAG, we caption images and tables and match them back to the retrieved text chunks before final generation, thereby maintaining a consistent candidate pool.
  • We conduct a systematic comparison of text-retrieval, image-retrieval, text–image fusion, and multimodal joint retrieval pipelines, analyzing which retrieval strategy performs best under different question types, evidence modalities, and document characteristics, providing practical guidance for choosing MM-RAG systems in real-world data settings.
@@ 2.1 Multimodal Retrieval-augmented generation (MM-RAG)
Recent advances in multimodal understanding highlight the importance of MM-RAG in reducing hallucinations. VLM2Vec (Jiang et al., 2024; Meng et al., 2025) demonstrated that instruction-tuning vision-language models significantly enhances their ability to produce robust embeddings, leading to strong performance across diverse text–image alignment tasks. Similarly, SeBe (Chen et al., 2025) adapts LLaVA-1.5 (Liu et al., 2024) by finetuning it into a retrieval-oriented embedding model, aligning user queries with external knowledge sources. GME (Zhang et al., 2024a) proposed a unified multimodal embedding model that is able to perform both text-to-image, image-to-text, and text-to-text retrieval. Uni-Retrieval (Jia et al., 2025) extends the paradigm by integrating VLMs with prompt-tuning strategies, enabling flexible handling of heterogeneous queries and modalities. Routing-based methods such as UniversalRAG (Yeo et al., 2025) and UniRAG (Sharifymoghaddam et al., 2025) introduce adaptive query routing mechanisms that dynamically select the most appropriate modality and level of granularity.
2.2 Visual Document Understanding and Evaluation
The challenge of document understanding with interleaved textual and visual components has recently prompted the development of specialized vision-based RAG pipelines (Yu et al., 2024a; Wang et al., 2025a; c) that directly take screenshots of documents as input. A notable example is ColPali (Faysse et al., 2024a), which leverages VLMs to jointly encode textual queries and visual documents with the MaxSim operations (Khattab & Zaharia, 2020). ViDoRAG (Wang et al., 2025a) introduces a multi-agent reasoning architecture designed for complex queries that require iterative cross-modal reasoning. In parallel, optimization-focused approaches such as VRAG (Wang et al., 2025c) apply reinforcement learning strategies, including GRPO-based Shao et al. (2024) training, to adapt VLMs for end-to-end document understanding. However, the comparisons with text-only baselines are not entirely fair, as most of these baselines exclude non-text modalities in response generation. Moreover, existing evaluations are conducted on datasets not designed for RAG. MMLongBench-Doc (Ma et al., 2024c) targets long-context multimodal document understanding, but its database is not highly relevant and thus unsuitable for retrieval tasks. REAL-MM (Wasserman et al., 2025) and VidoSeek (Wang et al., 2025b) are designed for MM-RAG, yet they lack cross-modality and multi-page evidence, limiting their ability to provide comprehensive and unified evaluation across RAG systems. The other benchmarks (Mathew et al., 2021; 2022; Zhu et al., 2022; Li et al., 2024) are typically limited to a single image or a single document page, covering narrow domains, under-representing modalities, or operating at limited scale with only a few queries or pages, as summarized in Table 1. To address these gaps, we introduce UniDoc-Bench, a benchmark tailored to practical RAG use cases.
@@ 3.1 Source Document Collection
We use PDFA (Montalvo & Wightman, 2024) as our data source, containing diverse formats (e.g., reports, slides, posters) and covering broad domains, but it lacks tags or labels. Therefore, our first step is data filtering to collect a high-quality database. We design a field scheme (see Appendix B.1) that captures key metadata, including domain, subdomain, language, modality (e.g., text, tables, figures), image quality (whether the resolution is clear), and text proportion. This allows us to standardize the data and build a high-quality cross-modality database. As shown in Figure 1 (c), we select $8$ domains based on differences across industries and define many subdomains within each, grouping similar documents. To ensure high inter-document similarity, we retain only documents from 353–5 related subdomains containing multiple modalities, yielding on average $8,000$ pages per domain. The final dataset spans Legal, Commerce and Manufacturing, Education, Energy, Construction, Finance, Healthcare, and CRM, with detailed subdomain descriptions in Appendix B.2.
Figure 2: Data Construction pipeline. (a) We filter and tag PDFA documents to curate a high-quality database of 7070k pages spanning 88 domains. (b) We parse documents into text, figures, and tables, then synthesize initial QA pairs covering four question types and three modalities using adapted templates. (c) We ground answers in supporting evidence, refine questions for human-intent and self-containment, and verify responses for factuality and completeness, yielding $1,600$ QA pairs. To ensure quality, $20\%$ of the dataset is validated by three independent human annotators.
@@ 3.2.1 Evidence Collection
PDF Parsing. We first parse our curated PDF document database [^2] https://unstructured.io/ by extracting text chunks, tables, and figures, with the latter two stored separately as image files. Within the parsed text chunk, each image and table is replaced with a unique placeholder tag (e.g., <<fig-XXX>> or <<tab-XXX>>), along with its corresponding caption and parsed content to fully represent interleaved multimodal content. An example of this parsing process is provided in Appendix B.3.
Chunks Grouping. To support multimodal evidence QA, we construct a knowledge graph ($\mathcal{G}_{i}$(ExplodingGradients, 2024; Peng et al., 2024) over the parsed chunks for domain $i$, where nodes ($N_{i}=\{n_{i1},n_{i2},$...\}) represent chunks and edges ($E_{i}$) denote overlapping entities (e.g., “AI Agent Platform”). Chunks across three modalities (text, tables, figures), from within or across documents, are linked to form ground-truth evidence, which are then used for QA synthesis in the next step.
@@ 3.2.2 Question and Answer Generation
Template Choice. First, we ensure the synthesized questions are diverse and span multiple categories, since focusing on a single category or using only the same few-shot example questions can introduce bias and limit the comprehensiveness of RAG evaluation. We designed four RAG question types: 1) factual retrieval, 2) comparison, 3) summarization, and 4) logical reasoning. For each question type and document domain, we design 10101515 general templates (see Appendix B.4). We then sample linked chunks ($n_{ij},e_{ij},n_{ik}$) and prompt the LLM to select 1133 templates ($T_{ij}$) that best match the provided chunks and are most likely to produce QA pairs that humans would naturally ask, thereby improving both the diversity and coverage of the questions.
Evidence Grounding. To ensure comprehensive evaluation of MM-RAG, we design four answer types with distinct evidence requirements, each supported by specialized prompts:
@@
  • Image-only: The question requires information exclusively from an image, such as numerical values shown only in a figure, thereby testing the system’s ability to interpret visual content.
  • Image-plus-text: Answering the question requires integrating information from both text and images, testing the model’s ability to reason across modalities.
@@
Rewriting. To ensure that questions are self-contained and reflect realistic human intent, we refine the initially synthesized QA pairs. In the first stage, many synthesized questions follow a long-context QA style and may include vague references such as “in this report” or “in Figure 8.” To make them suitable for RAG evaluation, we rewrite these questions to ensure they are self-contained and understandable without external context (Appendix B.6). Additionally, many QA pairs are grounded in images, leading to VQA-style questions (e.g., “How many logos are in Apple Inc.’s 2023 report?”). Such questions do not reflect natural human queries in a RAG context, so we filter and rewrite them to better align with realistic human intent. To ensure comprehensive evaluation, ground-truth answers must be complete and diverse. In the final step, we revise answers to cover all relevant aspects of their corresponding questions (see Appendix B.7).
Deduplication and Balance. Additionally, we remove duplicated question–answer pairs that are highly similar in the question or the answer (similarity >0.75>0.75333https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 ) to maintain dataset quality and diversity. We also rebalance the dataset by question type and answer type to provide a fair evaluation.
Dataset Statistics. Based on the above stages, we construct an evaluation benchmark consisting of $200$ QA pairs for each category, in total $1600$ QAs as described in Section 3.1. Within each set of $200$ QA pairs, we maintain an equal distribution of $50$ text-only, image-only, text-plus-image, and table-only questions. In total, the dataset contains $800$ single-modality and $800$ multi-modality questions. On average, each question requires $2.15$ evidence items (text chunks, images, or tables) for a complete answer, highlighting the need for RAG systems to retrieve multiple pieces of evidence. We further ensure a balanced distribution across the four main question types: factual retrieval, summarization, comparison, and logical reasoning. More details can be found in Figure 1(b).

### v3

@@ Abstract
Multimodal retrieval-augmented Generation (MM-RAG) is a key approach for applying large language models and agents to real-world knowledge bases, yet current evaluations are fragmented—focusing on either text or images in isolation, or simplified multimodal setup, failing to capture document-centric multimodal use cases. In this paper, we introduce UniDoc-Bench, the first large-scale, realistic benchmark for MM-RAG built from 7070k real-world PDF pages across $8$ domains. Our pipeline extracts and links evidence from text, tables, and figures, then generates $1,600$ multimodal QA pairs spanning factual retrieval, comparison, summarization, and logical reasoning queries. To ensure reliability, all of QA pairs are validated and rewritten by multiple human annotators and expert adjudication. UniDoc-Bench supports apples-to-apples comparison across four paradigms — 1) text-only, 2) image-only, 3) multimodal text–image fusion and 4) multimodal joint retrieval — under a unified protocol with standardized candidate pools, prompts, and evaluation metrics. UniDoc-Bench can also be used to evaluate Visual Question Answering tasks. Our experiments show that multimodal text–image fusion RAG systems outperform both unimodal and jointly multimodal embedding–based retrieval, indicating that neither text nor images alone are sufficient and that current multimodal embeddings remain inadequate. Beyond benchmarking, our analysis reveals when and how visual context complements textual evidence, uncovers systematic failure modes, and offers actionable guidance for developing more robust MM-RAG pipelines.
UniDoc-Bench: A Unified Benchmark for Document-Centric Multimodal RAG
Xiangyu Pengthanks: Equal contribution.    Can Qin11footnotemark: 1    Zeyuan Chen    Ran Xu    Caiming Xiong    Chien-Sheng Wu Salesforce AI Research {becky.peng, cqin, wu.jason}@salesforce.com
@@ 1 Introduction
Retrieval-augmented generation (RAG) has become a widely used approach for applying large language models (LLMs) and agents to real-world knowledge bases (gao2023retrieval; fan2024survey). The dominant text-only pipeline applies Optical Character Recognition (OCR) (li2022pp; xue2024xgen; poznanski2025olmocr) to flatten document pages into text, indexes them as chunks, retrieves top-k text passages, and feeds them to a generator. However, many answers depend on information embedded in figures, charts, tables, and complex layouts, where OCR often discards crucial spatial and visual semantics (e.g., map, axes, bar lengths, color encodings) (ma2024unifying; faysse2024colpali). These limitations have driven the rapid development of multimodal RAG (MM-RAG), which embeds documents across modalities (text, tables, and images) and retrieves and reasons over them jointly, emerging as a key paradigm for document intelligence.
@@
Current MM-RAG evaluation benchmarks exhibit substantial limitations, as summarized in Table 1. Many are restricted to a single image or a single document page as reference (mathew2021docvqa; mathew2022infographicvqa; zhu2022towards; li2024multimodal; ma2024mmlongbench), cover narrow domains mathew2021docvqa; mathew2022infographicvqa; zhu2022towards; li2024multimodal, under-represent modalities (li2024multimodal; mathew2022infographicvqa), operate at limited scale (few queries/pages)  (ma2024mmlongbench; wang2025vidorag) or lack a highly relevant database for RAG evaluation (ma2024mmlongbench). These gaps hinder fair and comprehensive comparison across methods. Moreover, debatable claims have emerged — such as that “image retrieval is all you need" (faysse2024colpali; su2025thinking) or that multimodal retrieval is inherently superior (zhang2024gme; yu2024visrag)— without enough fair and unified evaluation. In response, we introduce UniDoc-Bench, a human-verified benchmark spanning $8$ domains and covering text, chart, and table content, explicitly designed for cross-modality grounding with examples shown in Figure 1. Crucially, UniDoc-Bench enables apples-to-apples evaluation of text-retrieval, image-retrieval, multimodal text-image-fusion retrieval, and multimodal joint retrieval pipelines using highly relevant large document database and multi-type, cross-modality-grounding queries under a unified protocol. This setup provides an unbiased view of when multimodal retrieval offers advantages beyond single modalities. In practice, UniDoc-Bench quantifies multimodal gains, guides system design choices, and accelerates the development of effective MM-RAG systems for real-world document intelligence.
We curate a high-quality multimodal RAG evaluation benchmark by designing and applying a classification-based filtering scheme to unlabeled, real-world PDF documents (PDFA (montalvo_wightman_2024_pdfa_eng_wds)), yielding 7070k highly relevant pages across eight widely used domains —Finance, Legal, Healthcare, Commerce and Manufacturing, CRM, Energy, Education, and Construction—containing rich cross-modality content, including text, tables, and images. We construct a knowledge graph that links cross-modality contents across documents via overlapping entities, and leverage these connections to synthesize 1,600 QA pairs spanning four question types: factual retrieval, comparison, summarization, and logical reasoning, enabling multi-modality grounding and reflecting realistic retrieval scenarios. To ensure quality, all of the QA pairs are evaluated and rewritten by three independent annotators for faithfulness, completeness, self-containment, human intent, and evidence usability, with disagreements resolved through expert adjudication. Figure 2 illustrates the full pipeline from PDF segmentation to dataset creation and evaluation.
@@
Table 1: Comparison of existing document QA datasets with UniDoc-Bench.
Benchmarks Domain Evidence # Queries # Pages RAG Unified Multiple Human
of Doc Suitable Evaluation Reference Verif
ArxivQA (li2024multimodal) single [Uncaptioned image] 100k -
TAT-DQA (zhu2022towards) single [Uncaptioned image]  [Uncaptioned image] 17k 3k
InfoVQA (mathew2022infographicvqa) single [Uncaptioned image] 6k -
DocVQA (mathew2021docvqa) single [Uncaptioned image]  [Uncaptioned image] 11k -
MMLONG (ma2024mmlongbench) multiple [Uncaptioned image]  [Uncaptioned image][Uncaptioned image] 1.1k 5k
REALMM (wasserman2025real) multiple [Uncaptioned image]  [Uncaptioned image][Uncaptioned image] 5k 8k
ViDoSeek (wang2025vidorag) multiple [Uncaptioned image]  [Uncaptioned image][Uncaptioned image] 1.2k 10k
@@
  • Unified Evaluation: Apples-to-apples comparison across different baseline RAG systems.
  • Multiple Reference: Supports multi-hop, multi-modality, multi-source grounding.
  • Human Verif: Introduce human experts to review and verify the correctness and quality of all the QA pairs, or to annotate the entire dataset.
We make the following contributions:
  • We introduce a new multimodal RAG benchmark built from real-world PDF documents, comprising 7070k pages across $8$ domains, with $1,600$ human-verified QA pairs referencing text, figures, and tables, spanning $4$ question types.
  • We present a high-quality data synthesizing pipeline for creating MM-RAG evaluation datasets, designed to be compatible with any document database.
  • We propose a fair and reproducible evaluation framework by fixing candidate pools across modalities and measuring retrieval effectiveness, answer faithfulness, and completeness end-to-end across different RAG systems.
  • We compare text retrieval, image retrieval, text–image fusion, and multimodal joint retrieval pipelines, evaluating which strategy performs best across question types, evidence modalities, and document characteristics. We also show UniDoc-Bench’s use for evaluating Visual Question Answering (VQA) tasks, highlighting its versatility for MM-RAG research.
@@ 2.1 Multimodal Retrieval-augmented Generation (MM-RAG)
Recent advances in multimodal understanding underscore the importance of MM-RAG for reducing hallucinations. VLM2Vec (vlm2vec; meng2025vlm2vec) shows that instruction-tuning vision-language models improves embeddings for robust text–image alignment. SeBe (chen-etal-2025-seeing) adapts LLaVA-1.5 (liu2024improved) into a retrieval-oriented model that aligns user queries with external knowledge. GME (gme) proposes a unified multimodal embedding capable of text-to-image, image-to-text, and text-to-text retrieval. Uni-Retrieval (jia2025uni) combines VLMs with prompt-tuning to flexibly handle heterogeneous queries and modalities. Routing-based methods like UniversalRAG (yeo2025universalrag) and UniRAG (sharifymoghaddam2025unirag) use adaptive query routing to select the best modality and level of granularity.
2.2 Visual Document Evaluation
Document understanding with interleaved text and visuals has led to specialized vision-based RAG pipelines (visrag; vidorag; wang2025vrag) that process document screenshots directly. For example, ColPali (faysse2024colpali) uses VLMs to jointly encode textual queries and visual documents via MaxSim (khattab2020colbert), while ViDoRAG (vidorag) employs multi-agent reasoning for iterative cross-modal queries. Optimization-focused methods like VRAG (wang2025vrag) use GRPO (shao2024deepseekmath), to adapt VLMs for end-to-end document understanding. However, comparisons with text-only baselines are often unfair, as these baselines ignore non-text modalities. Existing evaluations are also limited: MMLongBench-Doc (mmlongbench) covers long-context multimodal documents but is poorly suited for retrieval; REAL-MM (wasserman2025real) and VidoSeek (wang2025vidorag) lack cross-page and cross-modal evidence; other benchmarks (mathew2021docvqa; mathew2022infographicvqa; zhu2022towards; li2024multimodal) are narrow in scope, covering single images or pages, limited domains, or small scales (Table 1). To fill these gaps, we introduce UniDoc-Bench, a benchmark designed for practical MM-RAG use cases with multi-page, cross-modal evidence and scalable evaluation.
@@ 3.1 Source Document Collection
We use PDFA (montalvo_wightman_2024_pdfa_eng_wds) as our data source, containing diverse formats (e.g., reports, slides, posters) and covering broad domains, but it lacks tags or labels. Therefore, our first step is data filtering to collect a high-quality database. We design a field scheme (Appendix B.1) that captures key metadata, including domain, subdomain, language, modality (e.g., text, tables, figures), image quality (whether the resolution is clear), and text proportion. This allows us to standardize the data and build a high-quality cross-modality database. As shown in Figure 1 (c), we select $8$ domains across industries and define subdomains within each, grouping similar documents. To ensure high inter-document similarity, we retain only documents from 353-5 related subdomains containing multiple modalities, yielding on $\sim 8,000$ pages per domain. The final dataset spans Legal, Commerce and Manufacturing, Education, Energy, Construction, Finance, Healthcare, and CRM, with detailed subdomain descriptions in Appendix B.2.
Figure 2: Data Construction pipeline. (a) We filter and tag PDFA documents to curate a high-quality database of 7070k pages spanning 88 domains. (b) We parse documents into text, figures, and tables, then synthesize initial QA pairs covering four question types and three modalities using adapted templates. (c) We ground answers in supporting evidence, refine questions for human-intent and self-containment, and verify responses for factuality and completeness, yielding $1,600$ QA pairs. To ensure quality, the entire dataset is validated and rewritten by human annotators.
@@ 3.2.1 Evidence Collection
PDF Parsing. We parse our curated PDF document database [^1] https://unstructured.io/ by extracting text chunks, tables, and figures, with the latter two stored separately as image files. Within the parsed text chunk, each image and table is replaced with a unique placeholder tag (e.g., <<fig-XXX>> or <<tab-YYY>>), along with its corresponding caption and parsed content to fully represent interleaved multimodal content. An example is provided in Appendix B.3.
Chunks Grouping. To support multimodal evidence QA, we construct a knowledge graph ($\mathcal{G}_{i}$(ragas2024; peng2024unanswerability) over the parsed chunks for domain $i$, where nodes ($N_{i}=\{n_{i1},n_{i2},$...\}) represent chunks and edges ($E_{i}$) denote overlapping entities (e.g., “AI Agent Platform”). Chunks across three modalities (text, tables, figures), from within or across documents, are linked to form ground-truth evidence, which are then used for QA synthesis in the next step.
@@ 3.2.2 Question and Answer Generation
Template Choice. First, we ensure the synthesized questions are diverse and span multiple categories, since focusing on a single category or using only the same few-shot example questions can introduce bias and limit the comprehensiveness of RAG evaluation. We designed $4$ RAG question types: 1) factual retrieval, 2) comparison, 3) summarization, and 4) logical reasoning. For each type and database domain, we design 10101515 templates (Appendix B.4). We then sample linked chunks ($n_{ij},e_{ij},n_{ik}$) and prompt the LLM to select 1133 templates ($T_{ij}$) that best match the provided chunks and are most likely to produce QA pairs that humans would naturally ask, thereby improving both the diversity and coverage of the questions.
Evidence Grounding. To ensure comprehensive evaluation of MM-RAG, we design $4$ answer types with distinct evidence requirements, each supported by specialized prompts:
@@
  • Image-only: The question requires information exclusively from an image, such as numerical values shown only in a figure.
  • Image-plus-text: Answering the question requires both text and images, testing the model’s ability to reason across modalities.
@@
Rewriting. To ensure that questions are self-contained and reflect realistic human intent, we refine the initially synthesized QA pairs. In the first stage, many synthesized questions follow a long-context QA style and may include vague references such as “in this report” or “in Figure 8”. To make them suitable for RAG evaluation, we rewrite these questions to ensure they are self-contained and understandable without external context (Appendix B.6). Also, many QA pairs are grounded in images, leading to VQA-style questions (e.g., “How many logos are in Apple Inc.’s 2023 report?”), which do not reflect natural human queries in a RAG context, so we filter and rewrite them to better align with realistic human intent. To ensure comprehensive evaluation, ground-truth answers must be complete and diverse. In the final step, we revise answers to cover all relevant aspects of their corresponding questions (see Appendix B.7).